OpenStudy (anonymous):

Determine whether each function is even, odd, or neither: y=x+sin^2 x

OpenStudy (unklerhaukus):

and odd function plus and even function is and odd function

OpenStudy (anonymous):

this is neither because y = -x + (sin(-x))^2 = -x +(-sinx)^2 = -x + (sinx)^2 therefore neither